If the efficiency of a 60W light bulb is 9.7%, how much energy is wasted due to heat in 2 hours?
1
Expert's answer
2020-05-21T13:32:37-0400

In case of 100% efficiency the wasted energy would be (by definition of the electrical power):

"A_{100\%} = 60W\cdot 2hours = 60W\cdot 7200 sec = 432kJ"

But in case of 9.7% efficiency:

"A_{9.7\%} = A_{100\%}\cdot 0.097 = 432kJ\cdot 0.097 = 41.9kJ"


Answer. A = 41.9 kJ.